Find the time it would take for the general level of prices in the economy to double at an average annual inflation rate of 7.1​%.
Answer by addingup(3677)

Find time for an annually compounding rate of 7.1% to double:
.
t = ln(A/P)/ln(1 + r) = [ln(A) - ln(P)] / ln(1 + r)
t = ln(2/1)/ln(1 + 0.071)
= [ln(2) - ln(1)]/ln(1 + 0.071)
= 0.3/0.03 = 10
It would take 10 years.